A group is to be picked from 6 seniors, 4 juniors, and 2 sophomores. if the committe of 3 is to be formed what is the probability that 1 of each will be selected?
Answer by stanbon(75887)

A group is to be picked from 6 seniors, 4 juniors, and 2 sophomores. if the committee of 3 is to be formed what is the probability that 1 of each will be selected?
--------------------
# of ways to get 1 of each = 6*4*2 = 48
---
# of possible groups of 3 = 12C3 = (12*11*10)/(1*2*3) = 220
----
Probability of 1 of each = 48/220 = 21818
=======================
Cheers,
Stan H.